Churchill Downs stock is expected to grow at 1.4% indefinitely, and the last dividend was 5$. If the relevant risk return is 4.2%, what would you be willing to pay for the stock.
Answer #1

Ans $ 181.07

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
P0 = 5.07 / (4.2%- 1.4%)
P0 = 181.07
D1 = D0* (1 + g)
D1 = 5* (1 + 1.4%)
D1 = 5.0700